Currently, the two fastest trains are the Japanese Maglev and the French TGV. The sum of their fastest speds is 718.2 mph. IF the speed of the Magleve is 3.8 mph faster than the speed of the TGV, find the speeds of each.
Answer by mananth(16949)

Currently, the two fastest trains are the Japanese Maglev and the French TGV. The sum of their fastest speds is 718.2 mph. IF the speed of the Magleve is 3.8 mph faster than the speed of the TGV, find the speeds of each.
Let the speed of the TGV be x mph
Speed of Maglev = x+3.8 mph
(x+3.8 + x) = 718.2
2x+3.8 = 718.2
2x= 718.2-3.80
2x= 714.4
X= 357.2 = TGV
357.2+3.8= 361 – Maglev mph
mananth@hotmail.com